OpenStudy (lgbasallote):

is this a homogeneous function? \[exp(\frac{x}{y})?\] i think it is a homo function with degree 0

OpenStudy (lgbasallote):

@TuringTest

OpenStudy (lgbasallote):

lol i forgot i could just tag =))

OpenStudy (turingtest):

it's not a function without an = sign

OpenStudy (lgbasallote):

oh,,well then expression...is it a homogeneous expression?

OpenStudy (turingtest):

that has no meaning.... what makes it homogeneous is the =0 part with out knowing if it is =0 or not we can't know if it is homogeneous anything

OpenStudy (lgbasallote):

uhh i guess it is =0...it's just how it was written in the book

OpenStudy (turingtest):

I've never heard of a "homogeneous expression" I should look it up

OpenStudy (lgbasallote):

lol...well i guess the point is just if it can be expressed as \[\lambda ^n f(x,y)\]

OpenStudy (turingtest):

if it is\[\exp(\frac xy)=0\]then that is non-homogeneous... but it's not even e DE where's the derivative sorry, Ill be back in 5...

OpenStudy (turingtest):

NH because because e^0=1....

OpenStudy (lgbasallote):

okay..i can wait :D curious question...are all "exp" non-homogeneous?

OpenStudy (blockcolder):

I think what he means is if \(f(x,y)=e^{x/y}\) is homogeneous. Right?

OpenStudy (lgbasallote):

yep

OpenStudy (blockcolder):

\[f(tx,ty)=e^{tx/ty}=e^{x/y}=t^0f(x,y)\] So it is homogeneous of degree 0.

OpenStudy (lgbasallote):

that's what i thought too...

OpenStudy (lgbasallote):

i guess turing is too tired right now to think straight ^_^